# Latent Representation Learning of Multi-scale Thermophysics: Application to Dynamics in Shocked Porous Energetic Material

### v1

@@ 3.1 Architecture Design
    learning a nonlinear representation of the slow manifold of the system together with the most probable fluctuations of the dynamics nearby the slow manifold; and
@@
Let $d$ be the dimension of each latent field $Z_{i}$, $i=1,\cdots,k$ over the spatial domain $\mathbb{X}$, i.e., the number of channels of the corresponding numerical tensor. In general, the convolution kernels $\bar{K}_{\theta_{i}}$ and ${K^{\sigma}_{\theta^{\prime}_{i}}}$, in (3.4) and (3.5), can be considered as $d$-by-$1$ matrix-valued functions over $\mathbb{X}\times\mathbb{X}$, i.e.,
@@
respectively, where $\mathcal{M}_{d\times 1}(\mathbb{R})$ denotes the set of all $d$-by-$1$ real matrices. In the particular setup of convolutional neural networks, one considers the convolution kernels to be translation-invariant matrix-valued functions with finite support, i.e., a fixed-size sliding window which sweeps the spatial domain $\mathbb{X}$. The entries of the corresponding matrices form the set of trainable parameters $(\theta_{i},\theta^{\prime}_{i})$ of the corresponding models which is learned through the training process. The latent mean field $\bar{Z}_{i}(x)$ (or, similarly, the log-variance field $Z^{\sigma}_{i}(x)$) at each $x\in\mathbb{X}$ is a weighted sum of $d$-dimensional vectors ${\bar{K}_{\theta_{i}}(x,y)\cdot U_{i}(y)}$ which are result of action of the matrix $\bar{K}_{\theta_{i}}(x,y)$ on the $1$-dimensional vector $U_{i}(y)$, $\forall y\in\mathbb{X}$.
@@
We consider a convolutional neural network (schmidhuber2015deep, ; mallat2016understanding, ) (CNN)-based architecture for all the building blocks of our model, i.e., the variational encoder, decoder, and the autoregressive functionals in (3.12) and (3.14), in order to make our model to be transferable from micro- to meso-scale. For the results presented in this paper, we consider a VGG-Net (simonyan2014very, ) and a U-Net (ronneberger2015u, ) architecture for the autoencoder and the latent dynamics autoregressive functionals, respectively. It is a multi-level, multi-resolution architecture which captures the local correlation between micro-scale building blocks in a hierarchical manner. Details of the implementation are discussed in Appendix A.1. Source codes are available online at https://github.com/ShahabAzarfar/Multiscale-Latent-Dynamics.
@@ 3.2 Training
Let ${\mathbf{E}=\left({\bar{\mathbf{E}},{\mathbf{E}}^{\sigma}}\right):\mathbf{U}\mapsto(\bar{\mathbf{Z}},{\mathbf{Z}}^{\sigma})}$ and ${\mathbf{D}:\mathbf{Z}\mapsto\widehat{\mathbf{U}}}$ be the direct sum of the VAE-encoders and the VAE-decoders, respectively, corresponding to the fields $U_{i}$, $i=1,\cdots,k$. Considering (3.9), the autoregressive rollout of our model for $M$ time-steps can be written as
@@ 4 Results
The thermomechanics governing the evolution of physical fields over a meso-scale domain (i.e., size of $O(100\,\mu m)$) of shock-loaded porous EM is the result of complex interactions between micro-scale dynamics resulting from the collapse of individual pores (saurel1999multiphase, ). The full set of governing equations for the corresponding dynamics is given as a system of hyperbolic PDEs, which involves the conservation laws for mass, momentum, and energy together with appropriate constitutive models such as the equation of state and flow rules. It describes the temporal evolution of a set of coupled physical fields, including the density, temperature, and pressure scalar fields; the velocity vector field; and the stress and strain 2-tensor fields (see (sambasivan2013simulation, ; kapahi2013dynamics, ; rai2017high, ) for details). The direct numerical simulation (DNS) data used in this paper is generated by solving the above-mentioned full system of coupled PDEs. More details on the numerical simulation setup are provided in Appendix A.2. In this work, following (nguyen2023parc, ), we consider the dynamical system corresponding to the evolution of the temperature $T$, pressure $P$ and microstructural morphology $\mu$ fields which are more directly related to the evolution of hotspots to quantify energy localization at the meso-scale to determine the macro-scale shock sensitivity quantities of interest for EM (udaykumar2020unified, ). These three fields play key roles in so-called reactive burn models which are commonly used to drive shock-to-detonation calculations in production hydrocodes (mader2007numerical, ). Our goal is to learn a surrogate model for the coupled dynamics of these three fields over a meso-scale domain in EM microstructures.
We follow a meta-learning approach to leverage a relatively large dataset of inexpensive micro-scale simulations, corresponding to single-pore collapse, combined with a small dataset of meso-scale simulations for the above-mentioned learning process. The proposed meta-learning framework is based on the idea that the meso-scale dynamics can be considered as a correlated system of interacting micro-scale building blocks, where the constituents can be learned from the underlying micro-scale physics. In this approach, we start by learning a latent representation of the micro-scale dynamics from a relatively large dataset ($O(100)$ samples) of micro-scale single pore collapse simulations. These machine-learned latent representations are considered as the building blocks of a “first order approximation” of the meso-scale dynamics. In the next step, in order to learn the “higher order corrections” corresponding to the correlation between the micro-scale building blocks, we train our model over a relatively small ($<10$ samples) dataset of meso-scale simulations. We elaborate on the above-mentioned two steps of our proposed meta-learning framework in Section 4.1 and 4.2, respectively.
@@ 4.1 Learning a latent representation of micro-scale dynamics
We start by learning a nonlinear representation of the slow manifold of the system. To this end, we train three independent VAEs, associated with $T$, $P$, and $\mu$ fields, on the single-void data in order to construct latent representations of the form $(\bar{Z}_{T},Z^{\sigma}_{T})$, $(\bar{Z}_{P},Z^{\sigma}_{P})$ and $(\bar{Z}_{\mu},Z^{\sigma}_{\mu})$, respectively. A visual 3D-embedding of the slow manifold of temperature latent mean fields $\bar{Z}_{T}$, corresponding to the hotspot ignition and growth due to shock-induced single pore collapse in EM, is shown in Figure 2.333We used a nonlinear dimensionality reduction technique, called PHATE (moon2019visualizing, ), for visualizing the manifold embedding. It is based on the Diffusion Maps algorithm (coifman2006diffusion, ) which defines an intrinsic notion of distance between the data points. One can see that the latent mean fields $\bar{Z}_{T}$ are initially quite close to each other which reflects the fact that, starting from the constant initial condition $T=300\mathrm{K}$, the temperature fields of all sample simulations, corresponding to the collapse of voids with various sizes and shapes, remain quite similar before the arrival of the shock front at the void which initiates the void-collapse and formation of hotspots. However, in later time-steps, the corresponding slow manifold expands and bifurcates into several branches depending on the temperature distribution and growth of the resulting hotspot.
@@
[…]
@@ 4.2 Learning the correlation between micro-scale building blocks
[…]
@@
[…]
@@
The differentiator module aims to learn the infinitesimal evolution of the corresponding physical fields encapsulated into $\widetilde{\mathbf{F}}$ while the integrator module computes $\Delta\mathbf{U}$ between two consecutive time-steps $[t_{i},t_{i}+\Delta t]$. The integrator module is considered to be a classical numerical integrator, e.g., Runge-Kutta methods, together with a neural network learning the higher-order corrections. In addition, to better capture the advection-dominant phenomenon, the PARCv2 differentiator architecture involves an inductive bias which is motivated by the generalized advection-diffusion-reaction equation.
In order to investigate the performance of both models in the scarce meso-scale data regime, we use a dataset containing 13 instances of shock-initiated reaction simulations in meso-scale microstructures. We consider four training scenarios for each model, using 10, 5, 3 and 1 training samples, respectively. Afterwards, the performance of each model is evaluated on 3 test samples kept separately during the training. Following our two-step meta-learning approach, our model has been trained, initially, on 100 samples of inexpensive micro-scale single-pore collapse simulations. However, PARCv2 is trained from scratch only on the meso-scale data. Figure 9 compares the convergence rate of two models in terms of their training loss in the case of 10 and 3 meso-scale training samples, respectively. We note that pretraining our model on the micro-scale data allows it to start from a lower initial loss value and converge relatively quickly compared to PARCv2, which is trained from scratch using only the meso-scale simulations.
@@
Figure 10 (a) and (b) display the predicted evolution of the temperature field $T$ by our model and PARCv2, respectively, for a test sample in comparison with the ground-truth numerical simulation. The columns in the figures show the predicted temperature distribution at different instants of time (noted above the panels); the rows correspond to the above-mentioned four scenarios considered for training of each model. The first row shows the ground truth, and the second up to the fifth row illustrate the prediction of each model trained on 10, 5, 3, and 1 meso-scale samples, respectively. In terms of capturing the formation and growth of the hotspots, our model’s prediction outperforms PARCv2, especially in the last three scenarios with few training samples. Our model’s predicted temperature values for the hotspots approaches closer to the ground-truth as the number of training samples increases, but it mostly stays below the ground-truth temperatures. On the other hand, PARCv2 tends to predict fewer hotspots with overpredicted temperature values. The boundary of our model’s predicted hotspots is blurrier compared with PARCv2, i.e., it has less sharp gradients relative to the ambient field. Also, similar observations hold in comparing the predicted evolution of the pressure field $P$ by our model and PARCv2 as illustrated in Figure 11 (a) and (b), respectively. PARCv2 can better capture the delicate patterns of the shock-waves in the pressure field compared to the present model which can be due to the fact that PARCv2 learns the dynamics in the high-fidelity input space instead of in the VAE-compressed latent space. We extend the comparison between the two models’ performance in terms of quantitative metrics in the following Subsection.
Figure 10: Comparison between the predicted temperature field evolution by (a) our model and (b) PARCv2 in four training scenarios.
Figure 11: Comparison between the predicted pressure field evolution by (a) our model and (b) PARCv2 in four training scenarios.
@@ 4.4 Quantitative evaluation of prediction performances
[…]

### v2

@@ 2.1 Architecture Design
    learning a nonlinear representation of the slow manifold of the system together with the most probable fluctuations of the dynamics in the neighborhood of the slow manifold; and
@@
Let $d$ be the dimension of each latent field $Z_{i}$, $i=1,\cdots,k$ over the spatial domain $\mathbb{X}$, i.e., the number of channels of the corresponding numerical tensor. In general, the convolution kernels $\bar{K}_{\theta_{i}}$ and $K^{\sigma}_{\theta^{\prime}_{i}}$}, in (2.4) and (2.5), can be considered as $d$-by-11 matrix-valued functions over $\mathbb{X}\times\mathbb{X}$, i.e.,
@@
respectively, where $\mathcal{M}_{d\times 1}(\mathbb{R})$ denotes the set of all $d$-by-11 real matrices. In the particular setup of convolutional neural networks, one considers the convolution kernels to be translation-invariant matrix-valued functions with finite support, i.e., a fixed-size sliding window which sweeps the spatial domain $\mathbb{X}$. The entries of the corresponding matrices form the set of trainable parameters $(\theta_{i},\theta^{\prime}_{i})$ of the corresponding models which is learned through the training process. The latent mean field $\bar{Z}_{i}(x)$ (or, similarly, the log-variance field $Z^{\sigma}_{i}(x)$) at each $x\in\mathbb{X}$ is a weighted sum of $d$-dimensional vectors $\bar{K}_{\theta_{i}}(x,y)\cdot U_{i}(y)$} which are result of action of the matrix $\bar{K}_{\theta_{i}}(x,y)$ on the $1$-dimensional vector $U_{i}(y)$, $\forall y\in\mathbb{X}$.
@@
We consider a convolutional neural network (schmidhuber2015deep, ; mallat2016understanding, ) (CNN)-based architecture for all the building blocks of our model, i.e., the variational encoder, decoder, and the autoregressive functionals in (2.12) and (2.14), in order to make our model to be transferable from micro- to meso-scale. For the results presented in this paper, we consider a VGG-Net (simonyan2014very, ) and a U-Net (ronneberger2015u, ) architecture for the autoencoder and the latent dynamics autoregressive functionals, respectively. It is a multi-level, multi-resolution architecture which captures the local correlation between micro-scale building blocks in a hierarchical manner. Note that the choice of CNN-based architecture is due to its local connectivity which leads to learn spatially local patterns in the input fields, and its ability to learn non-local effects as the depth of the architecture increases. Alternative well-established architectures for learning local and non-local patterns are Fourier Neural Operator (FNO) (li2020fourier, ) and Vision Transformer (ViT) (dosovitskiy2020image, ). However, these two architectures are computationally more expensive and less data efficient compared to CNNs which makes their applicability to the present scarce data scenario to be challenging. Details of the implementation of our CNN-based model are discussed in Appendix A.1. Source codes are available online at https://github.com/ShahabAzarfar/Multiscale-Latent-Dynamics.
@@ 2.2 Training
Let $\mathbf{E}=\left({\bar{\mathbf{E}},{\mathbf{E}}^{\sigma}}\right):\mathbf{U}\mapsto(\bar{\mathbf{Z}},{\mathbf{Z}}^{\sigma})$} and $^{\mathbf{D}:\mathbf{Z}\mapsto\widehat{\mathbf{U}}}$ be the direct sum of the VAE-encoders and the VAE-decoders, respectively, corresponding to the fields $U_{i}$, $i=1,\cdots,k$. Considering (2.9), the autoregressive rollout of our model for $M$ time-steps can be written as
@@ 3 Results
The thermomechanics governing the evolution of physical fields over a meso-scale domain (i.e., size of $O(100\,\mu m)$) of shock-loaded porous EM is the result of complex interactions between micro-scale dynamics resulting from the collapse of individual pores (saurel1999multiphase, ). The full set of governing equations for the corresponding dynamics is given as a system of hyperbolic PDEs, which involves the conservation laws for mass, momentum, and energy together with appropriate constitutive models such as the equation of state and flow rules. It describes the temporal evolution of a set of coupled physical fields, including the density, temperature, and pressure scalar fields; the velocity vector field; and the stress and strain 2-tensor fields (see (sambasivan2013simulation, ; kapahi2013dynamics, ; rai2017high, ) for details). The direct numerical simulation (DNS) dataset used in this paper is generated by solving the above-mentioned full system of coupled PDEs. More details on the numerical simulation setup are provided in Appendix A.2. In this work, following (nguyen2023parc, ), we consider the dynamical system corresponding to the evolution of the temperature $T$, pressure $P$ and microstructural morphology $\mu$ fields which are more directly related to the evolution of hotspots to quantify energy localization at the meso-scale to determine the macro-scale shock sensitivity quantities of interest for EM (udaykumar2020unified, ). These three fields play key roles in so-called reactive burn models which are commonly used to drive shock-to-detonation calculations in production hydrocodes (mader2007numerical, ). Our goal is to learn a surrogate model for the coupled dynamics of these three fields over a meso-scale domain in EM microstructures.
We follow a meta-learning approach to leverage a relatively large dataset of inexpensive micro-scale simulations, corresponding to single-pore collapse, combined with a small dataset of meso-scale simulations for the above-mentioned learning process. The proposed meta-learning framework is based on the idea that the meso-scale dynamics can be considered as a correlated system of interacting micro-scale building blocks, where the constituents can be learned from the underlying micro-scale physics. In this approach, we start by learning a latent representation of the micro-scale dynamics from a relatively large dataset ($O(100)$ samples) of micro-scale single pore collapse simulations. These machine-learned latent representations are considered as the building blocks of a “first order approximation” of the meso-scale dynamics. In the next step, in order to learn the “higher order corrections” corresponding to the correlation between the micro-scale building blocks, we train our model over a relatively small (<10<10 samples) dataset of meso-scale simulations. We elaborate on the above-mentioned two steps of our proposed meta-learning framework in Section 3.1 and 3.2, respectively.
@@ 3.1 Learning a latent representation of micro-scale dynamics
We start by learning a nonlinear representation of the slow manifold of the system. To this end, we train three independent VAEs, associated with $T$, $P$, and $\mu$ fields, on the single-void data in order to construct latent representations of the form $(\bar{Z}_{T},Z^{\sigma}_{T})$, $(\bar{Z}_{P},Z^{\sigma}_{P})$ and $(\bar{Z}_{\mu},Z^{\sigma}_{\mu})$, respectively. The encoder corresponding to each physical field in our model compresses the input field into a latent field with 64 times lower spatial resolution and four channels, which leads to a total compression rate equal to $1/16$. In other words, if we interpret the value of the involved physical fields in each spatial grid (pixel) as a degree of freedom of the corresponding multiphysics system, the autoencoder of our model learns a latent manifold embedded into the input space whose dimension equals $1/16$-th of the input space dimensionality. This level of dimensionality reduction makes it possible for our autoregressive model to learn a reduced representation of the original micro-scale dynamics which captures the main patterns of the underlying dynamics while ignoring its detailed high-frequency part.
A visual 3D-embedding of the slow manifold of temperature latent mean fields $\bar{Z}_{T}$, corresponding to the hotspot ignition and growth due to shock-induced single pore collapse in EM, is shown in Figure 2.333We used a nonlinear dimensionality reduction technique, called PHATE (moon2019visualizing, ), for visualizing the manifold embedding. It is based on the Diffusion Maps algorithm (coifman2006diffusion, ) which defines an intrinsic notion of distance between the data points. One can see that the latent mean fields $\bar{Z}_{T}$ are initially quite close to each other which reflects the fact that, starting from the constant initial condition $T=300\mathrm{K}$, the temperature fields of all sample simulations, corresponding to the collapse of voids with various sizes and shapes, remain quite similar before the arrival of the shock front at the void which initiates the void-collapse and formation of hotspots. However, in later time-steps, the corresponding slow manifold expands and bifurcates into several branches depending on the temperature distribution and growth of the resulting hotspot.
@@
A direct numerical simulation of the evolution of the temperature $T$ and pressure $P$ fields generated by shock-induced single pore collapse followed by hotspot ignition and growth, is illustrated in the top rows of Figure 5 (a) and (b), respectively. Roughly speaking, the time-periods $t\in[0,\,1.03\,\mathrm{ns}]$ and $t\in[1.03\,\mathrm{ns},\,2.75\,\mathrm{ns}]$ represent the above-mentioned two phases of the underlying dynamics, i.e., the ignition and the hotspot growth phases, respectively. The second rows of these figures show the global dynamics of the predicted temperature $\hat{T}$ and pressure field $\hat{P}$, respectively, resulting from the long-term rollout of our autoregressive model, given the initial condition of $T,P,\mu$ at $t=0$. Recall that, in our full model, the predicted temperature field $\hat{T}$ (resp. predicted pressure field $\hat{P}$) is reconstructed by the temperature variational decoder $D_{T}$ (resp. pressure variational decoder $D_{P}$) from the evolved temperature latent mean field $\bar{Z}_{T}$ and the temperature latent log-variance field $Z^{\sigma}_{T}$ (resp. pressure latent mean field $\bar{Z}_{P}$ and the pressure latent log-variance field $Z^{\sigma}_{P}$).
@@ 3.2 Learning the correlation between micro-scale building blocks
corresponding to the evolution of the state vector $\mathbf{U}=(T,P,\mu)$, where the parameters $\psi\in\Psi$ represent the geometrical and morphological characteristics of the underlying domain $\mathbb{X}_{\mathrm{mes}}$ at t=0t=0.555In this work, we restrict ourselves to the case of a fixed energetic material. Hence, the variations between the chemical properties of various energetic material is beyond the scope of this work. It is known that, depending on the geometrical properties of the underlying field of voids, the patterns of the temperature $T$ and pressure $P$ fields and their evolution, during the hotspot ignition and growth, can be quite different. In other words, the global behavior of the dynamical system (3.3) can change considerably by varying the geometric parameters $\psi$. In the following, we assume that the parametric solution map $\psi\mapsto\mathbf{U}_{\mathrm{mes}}(t;\psi)$ is a well-defined mapping from a low-dimensional manifold in the parameter space $\Psi$ to an appropriate function space over the time-interval $0,\mathbb{T}]$. In addition, we suppose that the meso-scale microstructures are sampled from a probability distribution $\mathcal{P}(\psi)$ over the parameter space $\Psi$. Using the parametric solution map $\psi\mapsto\mathbf{U}_{\mathrm{mes}}(\cdot\,;\psi)$, one can push-forward $\mathcal{P}(\psi)$ to the corresponding function space in order to get a stochastic process ${\{\mathbf{U}_{\mathrm{mes}}(t)\}}_{t\in[0,\mathbb{T}]}$ over the state space of the dynamical system (3.3). We use our probabilistic model to learn the underlying dynamics corresponding to this stochastic process.
@@
$\mathbf{F}_{\mathrm{mes}}=\mathbf{F}_{\mathrm{mic}}+\text{``higher-order correction terms''}\,$ (3.8)
@@
The differentiator module aims to learn the infinitesimal evolution of the corresponding physical fields encapsulated into $\widetilde{\mathbf{F}}$ while the integrator module computes $\Delta\mathbf{U}$ between two consecutive time-steps $t_{i},t_{i}+\Delta t]$. The integrator module is considered to be a classical numerical integrator, e.g., Runge-Kutta methods, together with a neural network learning the higher-order corrections. In addition, to better capture the advection-dominant phenomenon, the PARCv2 differentiator architecture involves an inductive bias which is motivated by the generalized advection-diffusion-reaction equation.
In order to investigate the performance of both models in the scarce meso-scale data regime, we use a dataset containing 13 instances of shock-initiated reaction simulations in meso-scale microstructures. We consider four training scenarios for each model, using 10, 5, 3 and 1 training samples, respectively. Afterwards, the performance of each model is evaluated on 3 test samples kept separately during the training. Following our two-step meta-learning approach, our model has been trained, initially, on 100 samples of inexpensive micro-scale single-pore collapse simulations. However, PARCv2 is trained from scratch only on the meso-scale data. Figure 9 compares the convergence rate of two models in terms of their training loss in the case of 10 and 3 meso-scale training samples, respectively. It can be seen that pretraining our model on the micro-scale data allows it to start from a lower initial loss value and converge relatively quickly compared to PARCv2, which is trained from scratch using only the meso-scale simulations. Note that the PARCv2 training loss is closely related to the $\mathcal{L}_{\mathrm{multi-step}}$ loss term, given by (2.19), considered for training our model in the sense that they both compare the predicted fields pixel-wise in the input space. However, we consider two extra loss terms $\mathcal{L}_{\mathrm{consistency-mean}}$ and $\mathcal{L}_{\mathrm{consistency-var}}$, given by (2.20) and (2.21), respectively, in training our autoregressive model which compare the predicted dynamics over latent space with the encoded dynamics using pretrained variational encoders.
@@
Figure 10 (a) and (b) display the predicted evolution of the temperature field $T$ by our model and PARCv2, respectively, for a test sample in comparison with the ground-truth numerical simulation. The columns in the figures show the predicted temperature distribution at different instants of time (noted above the panels); the rows correspond to the above-mentioned four scenarios considered for training of each model. The first row shows the ground truth, and the second up to the fifth row illustrate the prediction of each model trained on 10, 5, 3, and 1 meso-scale samples, respectively. In terms of capturing the formation and growth of the hotspots, our model’s prediction outperforms PARCv2, especially in the last three scenarios with few training samples. Our model’s predicted temperature values for the hotspots approaches closer to the ground-truth as the number of training samples increases, but it mostly stays below the ground-truth temperatures. On the other hand, PARCv2 tends to predict fewer hotspots with overpredicted temperature values. The boundary of our model’s predicted hotspots is blurrier compared with PARCv2, i.e., it has less sharp gradients relative to the ambient field. Also, similar observations hold in comparing the predicted evolution of the pressure field $P$ by our model and PARCv2 as illustrated in Figure 11 (a) and (b), respectively. PARCv2 can better capture the delicate patterns of the shock-waves in the pressure field compared to the present model which can be due to the fact that PARCv2 learns the dynamics in the high-fidelity input space instead of in the VAE-compressed latent space. The comparison between the predicted evolution of the temperature $T$ and pressure $P$ fields by our model and PARCv2 in the case of other test samples are presented in Appendix A.3 which confirm similar observations to the above-mentioned test sample. We extend the comparison between the two models’ performance in terms of quantitative metrics in the following Subsection.
Figure 10: Test sample #1: Comparison between the predicted temperature field evolution by (a) our model and (b) PARCv2 in four training scenarios.
Figure 11: Test sample #1: Comparison between the predicted pressure field evolution by (a) our model and (b) PARCv2 in four training scenarios.
@@ 3.4 Quantitative evaluation of prediction performances
Figure 12 (a) and (b) illustrate the time evolution of the total hotspot area $A_{\mathrm{hs}}$ and the average hotspot temperature $\bar{T}_{\mathrm{hs}}$ together with their rate of change for our model and PARCv2, respectively, in comparison with the ground-truth (DNS). All of these quantities of interest are calculated based on the predicted evolution of the temperature field for test samples by the two models trained on 10, 5, and 3 meso-scale samples, respectively. In terms of total hotspot area $A_{\mathrm{hs}}$ and its rate of change $\dot{A}_{\mathrm{hs}}$, our model’s prediction shows closer agreement with the ground-truth while PARCv2 tends to underpredict these two QoI. In addition, the overall trend of our model’s predicted average hotspot temperature $\bar{T}_{\mathrm{hs}}$ is consistent with the ground-truth, although it is underpredicted by $\sim 350\,\mathrm{K}$. On the other hand, PARCv2’s prediction of $\bar{T}_{\mathrm{hs}}$ can be divided into two parts based on the first and the second half of the total time-interval. Roughly speaking, PARCv2 underpredicts $\bar{T}_{\mathrm{hs}}$ for time-period $t\in[0,7.5\,\mathrm{ns}]$ while overpredicts $\bar{T}_{\mathrm{hs}}$ for $t\in[7.5\,\mathrm{ns},15\,\mathrm{ns}]$. The sharp increase in PARCv2 predicted $\bar{T}_{\mathrm{hs}}$ around $t=7.5\,\mathrm{ns}$ leads to considerable deviation of its predicted ${\dot{\bar{T}}}_{\mathrm{hs}}$ from ground-truth in that region. Figure 13 shows the root mean squared error (RMSE) of the predicted quantities of interest by the two models trained in the aforementioned four scenarios. The RMSE of our model’s predicted total hotspot area $A_{\mathrm{hs}}$ and its rate of change $\dot{A}_{\mathrm{hs}}$ is less than $\%50$ of the corresponding RMSE values for PARCv2 in all the training scenarios. On the other hand, the RMSE of PARCv2 predicted average hotspot temperature $\bar{T}_{\mathrm{hs}}$ decreases much faster than our model by increasing the number of training samples, leading to lower RMSE at the 10-sample training scenario.